OpenStudy (anonymous):

Ugh can someone show me the work for this? The length of a rectangle is 9.7 cm more than 4 times the width. If the perimeter of the rectangle is 91.4 cm, what are its dimensions?

OpenStudy (anonymous):

L = 9.7 + 4w

OpenStudy (anonymous):

Do you know perimeter formula of rectangular, then plug it in!

OpenStudy (anonymous):

p=2l+2w?

OpenStudy (anonymous):

2 ( W + L ) = P -> W + L = P/2 => W +9.7 + 4w = 91.4 /2 @whadduptori Now can you finish it up!